### v2

@@ 4.3 Synthetic Imaging
Using a similar approach as employed for the MUSTANG-2 case, Figure 13 shows the input map, the filtered map, and a synthetic observation of an output from the Diagona hydrodynamical simulation. The input maps represent the intracluster medium observed via the SZ effect from a cluster of galaxies. As an input, we used two clusters at , with one exhibiting a more compact surface brightness distribution than the other, but both showing features due to cluster merger activities. The chosen redshift corresponds to a FoV which corresponds to approximately Mpc, roughly equivalent to . The two clusters, with halo masses of , are shown in that specific order in the first column of Figure 13.
@@
5 Discussion and Summary
@@
The main drivers for this work were threefold. The first motivation is that the broader (sub-)mm community requires an easy-to-use, computationally efficient, and versatile simulator applicable to a variety of science cases, telescope sites, and designs. Section 2 provided the overview on how maria is set up to fulfill this goal. Second, maria enables a fair comparison between (sub-)mm observations with hydrodynamical simulations and other input models, as illustrated in Section 3. Henceforth, we can validate observational tools with these simulations. In particular, since the field of resolved SZ studies is now becoming more established, thanks to, for example, the wider bandwidth and sensitivity upgrades to ALMA (Carpenter et al., 2023) and the increased sensitivity of large single dish facilities, comparisons with hydrodynamical simulations are necessary to understand the systematic uncertainties in SZ analyses. The third and final motivation was to establish a connection between telescope designs, such as the number of detectors, field of view, and mapping speed, and the scientific goals for future facilities like the 50 meter Atacama Large Aperture Submillimeter Telescope (AtLAST). This objective was addressed in Section 4. However, regarding the latter goal, we simplified the instrument’s capabilities to facilitate simulation. More comprehensive, large-scale forecasts will be presented in follow-up works, utilizing the AtLAST design’s full field of view, polarization, and multichroic capabilities. For comparison, the data rates from mock observations will be larger when simulating a 2 FoV monochromatic instrument when simply filling AtLAST’s focal plane. This highlights the fact that AtLAST, with instrumentation fully populating the field of view, will be firmly situated in the Big Data regime.
@@
As this work illustrates in general the success of maria, it also has its limitations. Cosmic Microwave Background experiments or other single-dish simulators used for, for instance, the TolTEC Camera (see, e.g., Bryan et al., 2018) often refer to the Time-Ordered Astrophysics Scalable Tools package (TOAST, Puglisi et al., 2021; Kisner et al., 2023) or BoloCalc (Hill et al., 2018) to study systematic errors induced by the beam and calibration uncertainties. Apart from correlated atmospheric noise, TOAST and BoloCalc can also simulate cross-talk between detectors and scan-synchronous signals due to ground pickup as well as simulate leakage from total intensity to polarized light (Hill et al., 2018; Dachlythra et al., 2024). These noise contributions are not implemented in maria, nor are, for instance, surface deformations of the dish caused by gravity and thermal fluctuations. With maria, we simplistically assume diffraction-limited beams and uncorrelated pink and white noise. However, as demonstrated in Section 3, the noise corruption model implemented effectively replicates the noise characteristics observed in MUSTANG-2 data. Moreover, it’s easily adaptable to other telescope configurations at different sites. Additionally, our implementation of the location-and-time-specific atmosphere in maria stands out for its uniqueness while maintaining computational efficiency in computing the full evolving atmospheric volume.
@@
Looking forward, we are working to increase the flexibility of maria as future versions will encompass the ability to observe in an interferometric mode where vector averaged phase delays can be computed as a function of baseline length. This could be relevant for ALMA long baseline and high-frequency campaigns done, for instance, since maria could function as an improved phase screen model required to reproduce the measured phase fluctuations (see e.g., Maud et al., 2023). Furthermore, future implementations of maria will be better equipped to simulate direct detection spectrometers as was done, for instance, in TiEMPO (Huijten et al., 2022). Currently, we advise creating “moment-0” maps with a small frequency width, while future versions will be able to generate mock observations to imitate observations using heterodyne and integrated field units with higher spectral resolution.
@@
In closing, we emphasize that this paper is meant to introduce a virtual (sub-)mm telescope simulation tool that is generic and applicable across several science fields, from observations of the Sun (M. Kirkaune et al., in prep) to the circumgalactic medium of galaxies (Schimek et al. 2024, Lee et al., in prep). Even though the synthetic observations in this work all show the intracluster medium as seen via the SZ effect, the tool is broadly applicable and will be featured in future work.
Acknowledgements.
@@
This project has received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No. 951815 (AtLAST). MUSTANG-2 is supported by the NSF award number 1615604, by the Mt. Cuba Astronomical Foundation, and the Green Bank Observatory. The Green Bank Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. T. Morris acknowledges the support of L. Page. T. Morris acknowledges the musical work of Alan J. Lerner and Frederick Loewe as inspiration for the naming of the maria code. This work has been supported by the French government, through the UCAJ.E.D.I. Investments in the Future project managed by the National Research Agency (ANR) with the reference number ANR-15-IDEX-01. L.D.M. is supported by the ERC-StG “ClustersXCosmo” grant agreement 716762. L.D.M. further acknowledges financial contribution from the agreement ASI-INAF n.2017-14-H.0.References
- Abazajian et al. (2016) Abazajian, K. N., Adshead, P., Ahmed, Z., et al. 2016, arXiv e-prints, arXiv:1610.02743
- Ade et al. (2019) Ade, P., Aguirre, J., Ahmed, Z., et al. 2019, J. Cosmology Astropart. Phys., 2019, 056
- Assémat et al. (2006) Assémat, F., Wilson, R. W., & Gendron, E. 2006, Optics express, 14, 988
- Biffi et al. (2011) Biffi, V., Dolag, K., Boehringer, H., & Lemson, G. 2011, PHOX: X-ray Photon Simulator, Astrophysics Source Code Library, record ascl:1112.004
- Bleem et al. (2020) Bleem, L. E., Bocquet, S., Stalder, B., et al. 2020, ApJS, 247, 25
- Bleem et al. (2015) Bleem, L. E., Stalder, B., de Haan, T., et al. 2015, ApJS, 216, 27
- Bryan et al. (2018) Bryan, S., Austermann, J., Ferrusca, D., et al. 2018, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 10708, Millimeter, Submillimeter, and Far-Infrared Detectors and Instrumentation for Astronomy IX, ed. J. Zmuidzinas & J.-R. Gao, 107080J
- Carlstrom et al. (2011) Carlstrom, J. E., Ade, P. A. R., Aird, K. A., et al. 2011, PASP, 123, 568
- Carpenter et al. (2023) Carpenter, J., Brogan, C., Iono, D., & Mroczkowski, T. 2023, in Physics and Chemistry of Star Formation: The Dynamical ISM Across Time and Spatial Scales, 304
- CASA Team et al. (2022) CASA Team, Bean, B., Bhatnagar, S., et al. 2022, PASP, 134, 114501
- CCAT-Prime Collaboration (2023) CCAT-Prime Collaboration. 2023, ApJS, 264, 7
- Conan et al. (2000) Conan, R., Borgnino, J., Ziad, A., & Martin, F. 2000, JOSA A, 17, 1807
- Consortini & Ronchi (1972) Consortini, A. & Ronchi, L. 1972, Applied Optics, 11, 1205
- Coulton et al. (2023) Coulton, W. R., Madhavacheril, M. S., Duivenvoorden, A. J., et al. 2023, arXiv e-prints, arXiv:2307.01258
- Dachlythra et al. (2024) Dachlythra, N., Duivenvoorden, A. J., Gudmundsson, J. E., et al. 2024, ApJ, 961, 138
- Devlin et al. (2009) Devlin, M. J., Ade, P. A. R., Aretxaga, I., et al. 2009, Nature, 458, 737
- Di Mascolo et al. (2023) Di Mascolo, L., Saro, A., Mroczkowski, T., et al. 2023, Nature, 615, 809
- Dicker et al. (2014) Dicker, S. R., Ade, P. A. R., Aguirre, J., et al. 2014, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 9153, Millimeter, Submillimeter, and Far-Infrared Detectors and Instrumentation for Astronomy VII, ed. W. S. Holland & J. Zmuidzinas, 91530J
- Dünner et al. (2013) Dünner, R., Hasselfield, M., Marriage, T. A., et al. 2013, ApJ, 762, 10
- Errard et al. (2015) Errard, J., Ade, P. A. R., Akiba, Y., et al. 2015, ApJ, 809, 63
- Frayer (2017) Frayer, D. T. 2017, arXiv e-prints, arXiv:1706.02726
- Frayer et al. (2019) Frayer, D. T., Maddalena, R. J., White, S., et al. 2019, Calibration of Argus and the 4mm Receiver on the GBT, Green Bank Telescope Memorandum 302, June 5, 2019, 21 pages
- Gardini et al. (2004) Gardini, A., Rasia, E., Mazzotta, P., et al. 2004, MNRAS, 351, 505
- Henderson et al. (2016) Henderson, S., Allison, R., Austermann, J., et al. 2016, Journal of Low Temperature Physics, 184, 772
- Hersbach et al. (2020) Hersbach, H., Bell, B., Berrisford, P., et al. 2020, QJRMS, 146:730, 1999
- Hill et al. (2018) Hill, C. A., Bruno, S. M. M., Simon, S. M., et al. 2018, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 10708, Millimeter, Submillimeter, and Far-Infrared Detectors and Instrumentation for Astronomy IX, ed. J. Zmuidzinas & J.-R. Gao, 1070842
- Huijten et al. (2022) Huijten, E., Roelvink, Y., Brackenhoff, S. A., et al. 2022, Journal of Astronomical Telescopes, Instruments, and Systems, 8, 028005
- Jia et al. (2015) Jia, P., Cai, D., Wang, D., & Basden, A. 2015, Monthly Notices of the Royal Astronomical Society, 447, 3467
- Kaercher & Baars (2000) Kaercher, H. J. & Baars, J. W. 2000, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 4015, Radio Telescopes, ed. H. R. Butcher, 155–168
- Kisner et al. (2023) Kisner, T., Keskitalo, R., Zonca, A., et al. 2023, TOAST: Time Ordered Astrophysics Scalable Tools, Astrophysics Source Code Library, record ascl:2307.022
- Klaassen et al. (2020) Klaassen, P. D., Mroczkowski, T. K., Cicone, C., et al. 2020, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 11445, Ground-based and Airborne Telescopes VIII, ed. H. K. Marshall, J. Spyromilio, & T. Usuda, 114452F
- Lay (1997) Lay, O. P. 1997, A&AS, 122
- MacCrann et al. (2023) MacCrann, N., Sherwin, B. D., Qu, F. J., et al. 2023, arXiv e-prints, arXiv:2304.05196
- Madhavacheril et al. (2023) Madhavacheril, M. S., Qu, F. J., Sherwin, B. D., et al. 2023, arXiv e-prints, arXiv:2304.05203
- Maud et al. (2023) Maud, L. T., Asaki, Y., Nagai, H., et al. 2023, ApJS, 267, 24
- Morris et al. (2022) Morris, T. W., Bustos, R., Calabrese, E., et al. 2022, Phys. Rev. D, 105, 042004
- Morris et al. (2024 in prep) Morris, T. W., Page, L. A., & ACT Collaboration. 2024 in prep, in prep.
- Mroczkowski et al. (2023) Mroczkowski, T., Cicone, C., Reichert, M., et al. 2023, in 2023 XXXVth General Assembly and Scientific Symposium of the International Union of Radio Science (URSI GASS), 1–4
- Mroczkowski et al. (2019) Mroczkowski, T., Nagai, D., Basu, K., et al. 2019, Space Sci. Rev., 215, 17
- Paine (2018) Paine, S. 2018, The am atmospheric model
- Pardo et al. (2001) Pardo, J., Cernicharo, J., & Serabyn, E. 2001, IEEE Transactions on Antennas and Propagation, 49, 1683
- Plunkett et al. (2023) Plunkett, A., Hacar, A., Moser-Fischer, L., et al. 2023, PASP, 135, 034501
- Puglisi et al. (2021) Puglisi, G., Keskitalo, R., Kisner, T., & Borrill, J. D. 2021, Research Notes of the American Astronomical Society, 5, 137
- Qu et al. (2023) Qu, F. J., Sherwin, B. D., Madhavacheril, M. S., et al. 2023, arXiv e-prints, arXiv:2304.05202
- Ramasawmy et al. (2022) Ramasawmy, J., Klaassen, P. D., Cicone, C., et al. 2022, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 12190, Millimeter, Submillimeter, and Far-Infrared Detectors and Instrumentation for Astronomy XI, ed. J. Zmuidzinas & J.-R. Gao, 1219007
- Rasia et al. (2015) Rasia, E., Borgani, S., Murante, G., et al. 2015, ApJ, 813, L17
- Rasia et al. (2008) Rasia, E., Mazzotta, P., Bourdin, H., et al. 2008, ApJ, 674, 728
- Romero et al. (2020) Romero, C. E., Sievers, J., Ghirardini, V., et al. 2020, ApJ, 891, 90
- Ruze (1966) Ruze, J. 1966, Proceedings of the IEEE, 54, 633
- Schimek et al. (2024) Schimek, A., Decataldo, D., Shen, S., et al. 2024, A&A, 682, A98
- Sehgal et al. (2019) Sehgal, N., Aiola, S., Akrami, Y., et al. 2019, in Bulletin of the American Astronomical Society, Vol. 51, 6
- Sunyaev & Zeldovich (1970) Sunyaev, R. A. & Zeldovich, Y. B. 1970, Ap&SS, 7, 3
- Sunyaev & Zeldovich (1972) Sunyaev, R. A. & Zeldovich, Y. B. 1972, Comments on Astrophysics and Space Physics, 4, 173
- Swetz et al. (2011) Swetz, D. S., Ade, P. A. R., Amiri, M., et al. 2011, The Astrophysical Journal Supplement Series, 194, 41
- Tatarski (1961) Tatarski, V. I. 1961, Wave Propagation in a Turbulent Medium (McGraw-Hill, NY)
- The CMB-HD Collaboration et al. (2022) The CMB-HD Collaboration, :, Aiola, S., et al. 2022, arXiv e-prints, arXiv:2203.05728
- Thornton et al. (2016) Thornton, R. J., Ade, P. A. R., Aiola, S., et al. 2016, The Astrophysical Journal Supplement Series, 227, 21
- Tokovinin (2002) Tokovinin, A. 2002, Publications of the Astronomical Society of the Pacific, 114, 1156
- Wells et al. (1981) Wells, D. C., Greisen, E. W., & Harten, R. H. 1981, A&AS, 44, 363
- White et al. (2022) White, E., Ghigo, F. D., Prestage, R. M., et al. 2022, A&A, 659, A113
- ZuHone & Hallman (2016) ZuHone, J. A. & Hallman, E. J. 2016, pyXSIM: Synthetic X-ray observations generator, Astrophysics Source Code Library, record ascl:1608.002
- ZuHone et al. (2023) ZuHone, J. A., Vikhlinin, A., Tremblay, G. R., et al. 2023, SOXS: Simulated Observations of X-ray Sources, Astrophysics Source Code Library, record ascl:2301.024

### v3

@@ 4.3 Synthetic Imaging
Using a similar approach as employed for the MUSTANG-2 case, Figure 14 shows the input map, the filtered map, and a synthetic observation of an output from the Dianoga hydrodynamical simulation (see also section 3.3). The input maps represent the intracluster medium observed via the SZ effect from a cluster of galaxies. As an input, we used two clusters at , with one exhibiting a more compact surface brightness distribution than the other, but both showing features due to cluster merger activities. The chosen redshift corresponds to a FoV which corresponds to approximately Mpc, roughly equivalent to . The two clusters, with halo masses of , are shown in that specific order in the first column of Figure 14.
@@
5 Summary and Future Outlook
5.1 Summary
@@
The main drivers for this work were threefold. The first motivation is that the broader (sub-)mm community requires an easy-to-use, computationally efficient, and versatile simulator applicable to a variety of science cases, telescope sites, and designs. Section 2 provided the overview on how maria is set up to fulfill this goal. Second, maria enables a fair comparison between (sub-)mm observations with hydrodynamical simulations and other input models, as illustrated in Section 3. Henceforth, we can validate observational tools with these simulations. In particular, since the field of resolved SZ studies is now becoming more established, thanks to, for example, the wider bandwidth and sensitivity upgrades to ALMA (Carpenter et al., 2023) and the increased sensitivity of large single dish facilities, comparisons with hydrodynamical simulations are necessary to understand the systematic uncertainties in SZ analyses. The third and final motivation was to establish a connection between telescope designs, such as the number of detectors, field of view, and mapping speed, and the scientific goals for future facilities like the 50 meter Atacama Large Aperture Submillimeter Telescope (AtLAST). This objective was addressed in Section 4. However, regarding the latter goal, we simplified the instrument’s capabilities to facilitate simulation. More comprehensive, large-scale forecasts will be presented in follow-up works, utilizing the AtLAST design’s full field of view, polarization, and multichroic capabilities. For comparison, the data rates from mock observations will be larger when simulating a 2∘ FoV monochromatic instrument when simply filling AtLAST’s focal plane. This highlights the fact that AtLAST, with instrumentation fully populating the field of view, will be firmly situated in the Big Data regime.
@@
As this work illustrates in general the success of maria, it also has its limitations. Cosmic Microwave Background experiments or other single-dish simulators used for, for instance, the TolTEC Camera (see, e.g., Bryan et al., 2018) often refer to the Time-Ordered Astrophysics Scalable Tools package (TOAST, Puglisi et al., 2021; Kisner et al., 2023) or BoloCalc (Hill et al., 2018) to study systematic errors induced by the beam and calibration uncertainties. Apart from correlated atmospheric noise, TOAST and BoloCalc can also simulate cross-talk between detectors and scan-synchronous signals due to ground pickup as well as simulate leakage from total intensity to polarized light (Hill et al., 2018; Dachlythra et al., 2024). These noise contributions are not implemented in maria, nor are, for instance, surface deformations of the dish caused by gravity and thermal fluctuations. With maria, we simplistically assume diffraction-limited beams and uncorrelated pink and white noise. However, as demonstrated in Section 3, the noise corruption model implemented effectively replicates the noise characteristics observed in MUSTANG-2 data. Moreover, it is easily adaptable to other telescope configurations at different sites. Additionally, our implementation of the location-and-time-specific atmosphere in maria stands out for its uniqueness while maintaining computational efficiency in computing the full evolving atmospheric volume.
5.2 Future outlook
In the forthcoming era of big data and the global shift towards more sustainable astronomy, there is an urgent need to develop green telescopes that prioritize energy efficiency in both hardware and software. As mentioned, facilities like AtLAST, along with SO (Ade et al., 2019), SPT-3G (Benson et al., 2014), CMB-S4 (Abazajian et al., 2016), SKA (Dewdney et al., 2009), and ngVLA (Murphy et al., 2018) will generate large data volumes and require substantial storage and supercomputing resources to process and calibrate the data. One straightforward method to reduce computation time and storage requirements is to discard data, such as visibilities, or to down-sample time streams (as already done with MUSTANG-2 observations). However, these approaches could impact the reproducibility of research and limit the effectiveness of tools designed to extract the maximum scientific value from datasets. The above considerations underscore the need for more efficient reconstruction methods. In the case of maria, we will make significant improvements to the code infrastructure, such as enhancing parallelization and GPU acceleration by utilizing dask (Rocklin, 2015) and jax (Bradbury et al., 2018), which can boost performance by one or two orders of magnitude. This should make maria computationally efficient enough to simulate AtLAST and upcoming CMB-survey facilities to their full capacities.
In general, we expect that the most significant software advancement in the processing of ground-based CMB data will be achieved through the use of Gaussian processes. Although machine learning implementations might have a higher energy cost than simpler analysis tools, such as the naive mapmaking used in this work, they might be more efficient for retrieving scientific information and distinguishing the astronomical source from noise components. Single-dish facilities face significant challenges in separating atmospheric and astronomical signals, especially at large scales. This difficulty arises due to the large dynamic range in amplitudes—approximately seven orders of magnitude—between the atmosphere and the astronomical signal, as well as the fact that they share common modes across detectors, thus plausibly confusing large-scale signals for a common mode measured in the atmosphere. Additionally, the nonlinearity caused by sky scanning with a non-constant velocity further complicates the transformation from time-domain data to map-domain data. Traditional mapmaking methods assume a stationary, flat sky and pointing matrix creating a linear, invertible operation from time streams to sky coordinates (see, Dünner et al., 2013, section 11 for a more thorough description of mapmaking). However, this assumption is not valid at, for instance, the turnaround of a back-and-forth scan, eventually leading to loss of signal on large scales (Naess & Louis, 2023). Advancing map-making techniques and the corresponding steps for denoising timestreams from millions of detectors are well-suited to applications in information field theory, such as those relying on the Numerical Information Field Theory tool (NIFTy; Selig et al. 2013, Enßlin 2019, Edenhofer et al. 2024). NIFTy provides a powerful framework for inferring the true signal, which is obscured by a response function. In this context, it involves sampling the true signal and convolving the resulting detector time series with contributions from the atmosphere, the astronomical signal, and a noise component. The component separation in maria enables NIFTy to learn how to separate different components of the time series while accounting for non-linearity in sky sampling and the evolving atmosphere. These developments are currently in progress and will be explored and discussed in future work (Würzinger et al., in prep.).
@@ 5.2 Future outlook
A final application of maria to discuss is its potential for studying phase delays caused by the ever-changing turbulent atmosphere. This is particularly relevant for long-baseline, high-frequency campaigns conducted with ALMA, as maria could serve as an improved phase screen model, adding realism to replicate the empirical findings on measured phase stabilities reported in Maud et al. (2023). These findings are significant not only for ALMA but also for the ngVLA and would enable proper comparison of the maria code and observations in the terahertz regime, something currently unattainable with most single-dish facilities. Future implementations of maria will also be better equipped to simulate direct detection spectrometers as was done, for instance, in TiEMPO (Huijten et al., 2022). Currently, we advise creating “moment-0” maps with a small frequency width, while future versions will be able to generate mock observations to imitate observations using heterodyne and integrated field units with higher spectral resolution.
@@
In closing, we emphasize that this paper is meant to introduce a virtual (sub-)mm telescope simulation tool that is generic and applicable across several science fields, from observations of the Sun (Kirkaune et al., in prep) to the circumgalactic medium of galaxies (see, e.g., Lee et al., 2024; Schimek et al., 2024). Even though the synthetic observations in this work all show the intracluster medium as seen via the SZ effect, the tool is broadly applicable and will be featured in future work.
Acknowledgements
@@
This project has received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No. 951815 (AtLAST). MUSTANG-2 is supported by the NSF award number 1615604, by the Mt. Cuba Astronomical Foundation, and the Green Bank Observatory. The Green Bank Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. T. Morris acknowledges the support of L. Page. T. Morris acknowledges the musical work of Alan J. Lerner and Frederick Loewe as inspiration for the naming of the maria code. This work has been supported by the French government, through the UCAJ.E.D.I. Investments in the Future project managed by the National Research Agency (ANR) with the reference number ANR-15-IDEX-01. L.D.M. is supported by the ERC-StG “ClustersXCosmo” grant agreement 716762. L.D.M. further acknowledges financial contribution from the agreement ASI-INAF n.2017-14-H.0. JW is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2094 – 390783311.